OpenStudy (meilendurcer):

Will give medal please help. The function f(x) is shown below: f(x) = (4)x + 5 The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (4)^x - 8 f(x) = (4)^x + 13 f(x) = (4)^x + 18 f(x) = (4)^x - 13

OpenStudy (shubhamsrg):

any guess ?

OpenStudy (meilendurcer):

B

OpenStudy (shubhamsrg):

Close guess but no. try again.

OpenStudy (meilendurcer):

if it's close then could it be D?

OpenStudy (anonymous):

\[f(x) = 4^x + 5\]When you shift a function right by 'a' units, you subtract 'a' from its x variable so you get \[f(x-a)\]Here 'a' is 13, so you get \[f(x-13)=4^{x-13}+5\]

OpenStudy (anonymous):

None of the answer choices match the answer. The answer choices seem to be for shifting the function up or down, not right or left.

OpenStudy (anonymous):

For example, if you shift the function up by 13 units you get \[f(x)=4^x+18\]which is choice C

OpenStudy (meilendurcer):

@navk should i go with C?

OpenStudy (anonymous):

Either A or C because choice A corresponds to shifting the function down by 13 units while choice C corresponds to shifting it up by 13 units. Check the question again?

OpenStudy (shubhamsrg):

you are shifting f(x), so it should be D

OpenStudy (anonymous):

@shubhamsrg Can you elaborate a bit more as to how it can be D?

OpenStudy (shubhamsrg):

I am sorry I mis-interpreted. I was wrong. It won't be D.